OpenStudy (anonymous):

Help please! my username is mathdummy for a reason!...Find constants A and B that make the equation true 2x-9/x^2- x- 6 = A/x-3 + B/x+2

OpenStudy (anonymous):

lol smoothmath you probably think im a true math idiot by now

OpenStudy (anonymous):

First off i would like you to factor the denominator of the left hand side

hero (hero):

^What he said

OpenStudy (anonymous):

can't i multiply to set the denominators equal first?

OpenStudy (lgbasallote):

i see nothing suspicious in asking questions from substitution to calculus and then to partial fractions..

OpenStudy (anonymous):

if you can see it go ahead but i think that might be complicating the scenario

hero (hero):

lol @lgbasallote

OpenStudy (anonymous):

look igbasallote idk what to tell ya..if you don't believe me or dont want to believe me fine..but trust me when i say im not getting on here and asking math questions for fun...im taking several online courses and this just happens to be whats thrown at me

hero (hero):

Interesting.

OpenStudy (anonymous):

after you factor, and then multiply the denominator, you should have something like this 2x - 9 = A(x+2) + B(x-3)

OpenStudy (anonymous):

Do you understand how i got to that step

OpenStudy (anonymous):

ya that makes sense to me...

OpenStudy (anonymous):

now, the next step will be to distribute the A and the B Show me what you get when you finish

OpenStudy (anonymous):

2x-9 = Ax + 2A + Bx - 3B.... 2x = Ax + Bx -9 = 2A - 3B 2 = A + B -9 = 2A - 3B 6 = 3A + 3B -9 = 2A - 3B +.....am i doing this right so far??

hero (hero):

Looks right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

At this point, 2 = A + B -9 = 2A - 3B we can do substitution. 2 = A + B, can be written as ; A = 2 - B

OpenStudy (anonymous):

then substitute A into the other equation ( -9 = 2A - 3B )

OpenStudy (anonymous):

try it it should be easy from here. First you will get B. And then you can solve for A

OpenStudy (anonymous):

Okay so for my final answer i got B = 13/5 and A = 3/5...look right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks so much for your help!!

OpenStudy (anonymous):

Anytime